You have been managing a $5 million portfolio that has a beta of 1.45 and a required rate of return of 12.425%. The current risk-free rate is 3%. Assume that you receive another $500,000. If you invest the money in a stock with a beta of 1.15, what will be the required return on your $5.5 million portfolio?

Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

Currently:

12.425=3+1.45*(market rate-3)

(12.425-3)=1.45*(market rate-3)

market rate=(12.425-3)/1.45+3

=9.5%

Hence required return for $500,000=3+1.15*(9.5-3)

=10.475%

Portfolio return=Respective return*Respective weight

=(5,000,000/5,500,000*12.425)+(500,000/5,500,000*10.475)

=12.25%(Approx)